# Vulnerability Disclosure or Notification? Best Practices for Reaching Stakeholders at Scale

### v1

@@ 6.2. Communication Channel
Choosing the most effective channels to perform disclosure is the key to reaching out to stakeholders (Wogalter et al., 1999). In the case of vulnerability disclosure to vendors, there is usually a vulnerability disclosure policy in place (Boucher and Anderson, 2022; Moura and Heidemann, 2023), possibly indicated using a security.txt (Foudil and Shafranovich, 2022). The vendor may also have outsourced this to third-party platforms such as HackerOne and Bugcrowd (Walshe and Simpson, 2020a), and having a legal agreement in disclosure (Macnish and van der Ham, 2020). If pre-assessment on the potential stakeholders or impact of a vulnerability is unclear, one can reach out to national CSIRTs to consult with a possible communication channel (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]). National CSIRTs may provide notification services, coordination support or platforms such as VINCE by CERT/CC (CERT/CC, [n. d.]) to efficiently identify and inform vendors (Forum of Incident Response and Security Teams (2020), FIRST; National Cyber Security Centre of The Netherlands (2018), NCSC-NL).
For vulnerability notification to end-users, finding the right communication channel is a big challenge. One can first consider selecting known stakeholders with clear contact information (Moura and Heidemann, 2023; Beitis and Vanhoef, [n. d.]), then seek the rest of the stakeholders’ contact information. However, one should also be mindful of the messaging prioritisation and scheduling if the response time from the initial list of stakeholders takes longer than expected (Moura and Heidemann, 2023). Moreover, one should be aware that third-party delegation to national CSIRTs may not be the most effective method, as the interest, notification coverage, and capacity limitation of national CSIRTs may not align with the stakeholders (Durumeric et al., 2014; Çetin et al., 2017a; Stock et al., 2018b).
We do not recommend using email to perform large-scale vulnerability notifications to end-users, even though this has been the most commonly used option. Setting up an email infrastructure for doing large-scale notifications is difficult, keeping in mind all of the spam prevention tools currently in use. This comes on top of the difficulty of actually finding contact information for systems on the Internet. The accuracy of the contact information presented on web pages is not guaranteed (Maass et al., 2021). Finding contact information for IP addresses is notoriously hard. Although WHOIS and RDAP are methods to retrieve the contact information, the current results of both are often not accurate (Fernandez et al., 2024). Moreover, the contact information available is often meant for abuse notifications, not for vulnerability notifications. While the ‘security.txt’ standard(Foudil and Shafranovich, 2022) works for websites, there is no such alternative for IP addresses.
@@
Depending on the affected parties, one can consider reaching out to key communities or platforms to support remediation (Boucher and Anderson, 2022; Moura and Heidemann, 2023). For large-scale notifications, the use of Shadowserver as a notification channel seems most promising (Beitis and Vanhoef, [n. d.]). Shadowserver can perform active scans and has trusting relationships with key stakeholders that support the notification infrastructure. This has the added advantage that end-users are not overwhelmed with notification campaigns from different researchers. Phone calls and posts are more of an option when stakeholders suggest doing so (Maass et al., 2021; Stock et al., 2018b).
@@ 6.3. Messaging Infrastructure
The messaging infrastructure will depend on the selected communication channels. Researchers should follow the indicated preference of vendors in using the contact method, usually email, or messaging platforms such as bug bounty or VINCE.
If an email infrastructure is still used, then we recommend that the sending email address be from a known domain name or with an organisation to increase the delivery rate (van der Horst, 2023; Stock et al., 2018b; Chen et al., 2024). One should also be aware of the implementation of their mail service; with the mail service outsourcing trend, there can be rate limits and extra policies to examine beforehand (Chen et al., 2024). Be aware that recipients may send automatic responses, divert to ticketing systems, request feedback forms (Çetin et al., 2017a; Durumeric et al., 2014; Chen et al., 2024; Stock et al., 2018b), or outsourced platforms with different disclosure policies (Boucher and Anderson, 2022; Van Hove et al., 2023). It should be noted that the effectiveness of large-scale email notifications leaves a lot to be desired, as can be seen in Table 2.
Aside from the mail service, one can also provide web pages to describe the intention of the disclosure, vulnerability information, remediation, and disclosure policy. This can help reduce the content in email messages and provide a static source for stakeholders to help track the issues in the long term (Vanhoef, [n. d.]; Çetin et al., 2017a; Stock et al., 2018b; Chen et al., 2024).
The scalability of the messaging infrastructure is the biggest challenge when performing large-scale vulnerability notifications. Although organisations such as CERT/CC with VINCE and national CSIRTs with vulnerability notification systems can help with large-scale disclosure, there is still no optimised channel and infrastructure for vulnerability notifications. Shadowserver seems to be a promising infrastructure for effectively reaching affected parties.
@@ 6.4. Disclosure Policy and Message
Composing a disclosure message is never trivial. Among stakeholders, a finder and affected parties can have different preferences and procedures for handling the messages (Walshe and Simpson, 2022b; van der Horst, 2023). In vulnerability disclosure, vendors with bounty programs or disclosure policies may provide clear instructions or at least contact information to initiate the disclosure (Walshe and Simpson, 2020a). In contrast, recipients of the vulnerability notification may not provide enough information on how they will handle a disclosure message. Network operators may prefer extensive information that includes more details and remediation (van der Horst, 2023; National Cyber Security Centre of The Netherlands (2018), NCSC-NL), given that the affected party follows a certain time constraint policy on remediation (Chen et al., 2024). Meanwhile, cloud providers or domain owners may forward the message to their clients with limited communications (Chen et al., 2024; Li et al., 2016). Eventually, one should be aware that not every stakeholder would accept longer messages or attachments regarding the mail server filter, ticketing systems, and spam filter (Çetin et al., 2017a; Stock et al., 2018b; Chen et al., 2024). As a result, we recommend that a finder ensure that the initial message remains brief and does not necessarily reveal every detail of the vulnerability in case of an information leak or legal action (van der Ham et al., 2023; van der Horst, 2023).
@@
In conversations with stakeholders, one should consider that large-scale vulnerability disclosure and notification may get manual or automatic responses from various communication systems in different languages. In most cases, we have seen English used in the responses; other languages are used as support (Durumeric et al., 2014; Li et al., 2016; Chen et al., 2024). Messages in different languages may increase the processing time and cause confusion, especially if disclosure policies are presented in non-native languages to a finder or a receiving party. Besides, when using stakeholders’ communication systems, a finder may need to put extra effort into reformulating the messages based on the limitations of the disclosure form or editors in the provided text system (Chen et al., 2024).
All things considered, we recommend that a finder and support organisations establish an outgoing disclosure policy (van der Ham et al., 2023; Reidsma et al., 2023; National Cyber Security Centre of The Netherlands (2018), NCSC-NL), which includes legal terms, disclosure schedules, message templates, or exemptions, as implemented in (Stivala et al., 2024; Van Hove et al., 2023; Chen et al., 2024). This can let recipients understand the disclosure procedure from a finder and protect the finder from unwanted behaviours such as legal actions (Macnish and van der Ham, 2020) and public criticism (Van Hove et al., 2023). Moreover, we recommend that the stakeholders include a disclosure policy(CERT/CC, 2024; National Cyber Security Centre of The Netherlands (2018), NCSC-NL; van der Ham et al., 2023) or security.txt(Foudil and Shafranovich, 2022) in their disclosure programs or responses. This helps a finder to initiate the disclosure and notification operations with the stakeholders with better preparation. As we observed in the community and selected work, disclosure policies and equivalent documentation are not widely implemented, and this has resulted in both finders and receiving parties having inefficient communication.
@@
Last but not least, one should be aware of the stance of stakeholders on disclosure or notification best practices. Not every stakeholder may want to adopt the best practices for different reasons. This has been documented in two vulnerability disclosure operations (Vanhoef, [n. d.]; Van Hove et al., 2023) with vendors refusing to participate in the remediation schedule despite having a national CSIRT as coordinator. Still, we encourage stakeholders to adopt best practices and participate in disclosure or notification operations.
@@ 6.5. Post-Review
Measuring the impact after vulnerability disclosure and notification can help reflect on this process and improve best practices. If a vulnerability or a security issue is traceable by performing network scanning, measuring with stakeholders, or communicating with end-users, it can be valuable to record the remaining number of vulnerable systems after disclosure or notification. Additionally, providing a timeline as figure (Boucher and Anderson, 2022; Van Hove et al., 2023; Moura and Heidemann, 2023; Durumeric et al., 2014; Stock et al., 2018b, 2016a) or as text (Chen et al., 2024; Stivala et al., 2024) for each disclosure stage can help understand the operation development and its impact over time. These aspects have been covered well in our selected work, revealing the tradeoffs and considerations in different scenarios. Aside from monitoring vulnerable systems and messaging infrastructure, what the finders encountered and stressed during their operation is also worth recording. As revealed in nearly all reviewed publications, handling disclosure or notification procedures can be a nerve-racking trial. It is crucial to have evolving best practices to help a finder prepare for the challenges and ease the stress during disclosure operations. With the fast-changing nature of the network landscape and the growing number of vulnerabilities, it is important to have best practices up to date and mitigate security issues in time.
We have seen publications focusing on the exploits, attacks, and network traffic before and after the public disclosure. However, the work focusing on large-scale vulnerability disclosure and notification is relatively less seen in the community, which is also mentioned in (Chen et al., 2024; Moura and Heidemann, 2023; Stock et al., 2018b). Further, we have observed that not every finder has the capacity and prior experience to conduct disclosure or notification operations and finish the documentation within the academic publication period. Nevertheless, we have seen researchers in (Chen et al., 2024; Stivala et al., 2024; Beitis and Vanhoef, [n. d.]) could learn from the prior best practices and contribute to large-scale notification practices. Therefore, we encourage researchers, ethical hackers, practitioners, and operators to document their disclosure or notification operations with their own experiences, considerations, and outcomes. In the selected work, the tracking period can range from weeks (Durumeric et al., 2014), months (Çetin et al., 2019b; Chen et al., 2024), or more than a year (Stivala et al., 2024) if the situation permits. It could be a webpage (Vanhoef and Piessens, 2017a; Boucher and Anderson, [n. d.]) and followup publications (Durumeric et al., 2014; Stivala et al., 2024; Boucher and Anderson, 2022; Stock et al., 2018b; Vanhoef and Piessens, 2018b) to present disclosure or notification updates.

### v2

@@ 5.2. Communication Channel
Choosing the most effective channels to perform disclosure is the key to reaching out to stakeholders (Wogalter et al., 1999). We provide our suggestions on communication channels for large-scale vulnerability disclosure and notification separately.
In the case of vulnerability disclosure to vendors, there are already comprehensive guidelines and support mechanisms in place to help finders inform the receiving stakeholders:
  • A finder can seek vulnerability disclosure policies or programs of stakeholders (Boucher and Anderson, 2022; Moura and Heidemann, 2023), which are possibly indicated by stakeholders providing a security.txt (Foudil and Shafranovich, 2022).
  • A finder can also look for bug bounty programs or vulnerability platforms to issue the vulnerability report. However, stakeholders may also have outsourced such channels to third-party platforms such as HackerOne and Bugcrowd (Walshe and Simpson, 2020a), and having a legal agreement in disclosure (Macnish and van der Ham, 2020).
@@
  • A finder can reach out to national CSIRTs to consult on a possible communication channel (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]), if the pre-assessment of potential stakeholders or the impact of a vulnerability is unclear. National CSIRTs may provide notification services, coordination support or platforms such as VINCE by CERT/CC (CERT/CC, [n. d.]) to efficiently identify and inform vendors (Forum of Incident Response and Security Teams (2020), FIRST; National Cyber Security Centre of The Netherlands (2018), NCSC-NL).
In the case of vulnerability notification to end-users, finding an effective and efficient communication channel still remains a significant challenge despite the current guidelines:
  • A finder can first consider selecting known stakeholders with clear contact information (Moura and Heidemann, 2023; Beitis and Vanhoef, [n. d.]), then seek the rest of the stakeholders’ contact information.
  • A finder should then be mindful of the messaging prioritisation and scheduling if the response time from the initial list of stakeholders takes longer than expected (Moura and Heidemann, 2023).
  • A finder should be aware that delegation to national CSIRTs mostly works but may not be effective in every scenario, as the interests, notification coverage, and capacity of different national CSIRTs or equivalent support organisations may vary (Durumeric et al., 2014; Çetin et al., 2017a; Stock et al., 2018b).
@@
  • A finder should consider reaching out to key communities or platforms that support remediation tracking (Boucher and Anderson, 2022; Moura and Heidemann, 2023), depending on the affected parties. Phone calls and posts are more of an option when stakeholders suggest doing so or the regimes have such practice (Maass et al., 2021; Stock et al., 2018b).
  • We do not recommend using email to perform large-scale vulnerability notifications to end-users, even though this has been the most commonly used option. Setting up an email infrastructure for doing large-scale notifications is difficult, keeping in mind all of the spam prevention tools currently in use.
  • We recommend the use of The Shadowserver Foundation as a notification channel, as implemented by Beitis and Vanhoef (Beitis and Vanhoef, [n. d.]). The organisation can perform active scans and has trusting relationships with key stakeholders that support the notification infrastructure. This has the added advantage that end-users are not overwhelmed with notification campaigns from different finders.
Email as a communication channel in large-scale notifications faces the additional challenge of finding contact information for systems on the Internet. The accuracy of the contact information presented on web pages is not guaranteed (Maass et al., 2021). Finding contact information for IP addresses is notoriously hard. Although WHOIS and RDAP are methods to retrieve the contact information, the current results of both are often not accurate (Fernandez et al., 2024). Moreover, the contact information available is often meant for abuse notifications, not for vulnerability notifications. While the ‘security.txt’ standard(Foudil and Shafranovich, 2022) works for websites, there is no such alternative for IP addresses.
@@ 5.3. Messaging Infrastructure
The messaging infrastructure will depend on the selected communication channels. Finders should follow the indicated preference of vendors in using the contact method, usually email, or messaging platforms such as bug bounty programs, VINCE, or other vulnerability report platforms. It should be noted that the effectiveness of large-scale email notifications leaves a lot to be desired, as can be seen in Table 2. If an email infrastructure is still used:
  • We recommend that the sending email address be from a known domain name or with an organisation to increase the delivery rate (van der Horst, 2023; Stock et al., 2018b; Chen et al., 2024). One should also be aware of the implementation of their mail service; with the mail service outsourcing trend, there can be rate limits and extra policies to examine beforehand (Chen et al., 2024).
  • A finder should be aware that recipients may send automatic responses, divert to ticketing systems, internal communication or management systems, request feedback forms (Çetin et al., 2017a; Durumeric et al., 2014; Chen et al., 2024; Stock et al., 2018b), or outsourced platforms with different disclosure policies (Boucher and Anderson, 2022; Van Hove et al., 2023).
The scalability of the messaging infrastructure is the biggest challenge when performing large-scale vulnerability notifications. Although organisations such as CERT/CC with VINCE and national CSIRTs with vulnerability notification systems can help with large-scale disclosure, there is still no optimised channel and infrastructure for large-scale vulnerability notifications in the existing guidelines (Imp, [n. d.]). Nonetherless, aside from the email infrastructure:
  • A finder can provide web pages to describe the intention of the disclosure, vulnerability information, remediation, and disclosure policy. This can help reduce the content in email messages and provide a static source for stakeholders to help track the issues in the long term (Vanhoef, [n. d.]; Çetin et al., 2017a; Stock et al., 2018b; Chen et al., 2024).
  • We recommend checking with a support organisation or partner for messaging infrastructure. National CSIRTs, PSIRTs, ISPs, and more stakeholders have been improving external or internal notification mechanisms (Inf, [n. d.]; Çetin et al., 2019b; Rodríguez et al., 2021).
  • We recommend The Shadowserver Foundation, with its established infrastructure and notification experiences, as an effective option for network scanning, identifying vulnerable systems, and reaching affected parties (Sha, [n. d.]).
In addition, it is worth noting that the adoption of bug bounty programs and vulnerability report platforms as accepted best practices among stakeholders has been growing over the years. However, while well-established stakeholders may have mature report-handling policies and support mechanisms to receive vulnerability reports at scale (Boucher and Anderson, 2022), other stakeholders may still lack the capacity and experience to present a clear disclosure policy or program in their public information or ticketing system (Chen et al., 2024). Although some regimes with laws and regulations (cra, [n. d.]; BOD, 2020; National Cyber Security Centre of The Netherlands (2018), NCSC-NL) that mandate stakeholders to take action upon notification have successful cases, such as (Maass et al., 2021), it is not guaranteed that stakeholders will initiate mitigation, particularly in multi-party notifications at the Internet-wide scale with diverse stakeholders involved, as presented in (van der Horst, 2023; Chen et al., 2024). Moreover, even vendors or other stakeholders may still struggle to notify their clients and ensure trust from the receiving end-users with correct contact information (Imp, [n. d.]). Eventually, it will be easier for finders to perform vulnerability disclosure and notification at scale once the adoption of the practices gets higher.
@@ 5.4. Disclosure Policy and Message
Composing a disclosure message is never trivial. Among stakeholders, a finder and affected parties can have different preferences and procedures for handling the messages (Walshe and Simpson, 2022b; van der Horst, 2023).
  • In vulnerability disclosure, stakeholders with bounty programs or disclosure policies may provide clear instructions or at least contact information to initiate the disclosure (Walshe and Simpson, 2020a). In contrast, recipients of the vulnerability notification may not provide enough information on how they will handle a disclosure message.
  • A finder should know that stakeholders like network operators may prefer extensive information that includes more details and remediation (van der Horst, 2023; National Cyber Security Centre of The Netherlands (2018), NCSC-NL), given that the affected party follows a certain time constraint policy on remediation (Chen et al., 2024). Meanwhile, stakeholders like cloud providers or domain owners may forward the message to their clients with limited communications (Chen et al., 2024; Li et al., 2016).
  • A finder should be aware that not every stakeholder would accept longer messages or attachments regarding the mail server filter, ticketing systems, and spam filter (Çetin et al., 2017a; Stock et al., 2018b; Chen et al., 2024).
  • We recommend that a finder ensure that the initial message remains brief and does not necessarily reveal every detail of the vulnerability in case of an information leak or legal action (van der Ham et al., 2023; van der Horst, 2023).
In conversations with stakeholders after initiating vulnerability disclosure or notification.
@@
  • A finder should consider that large-scale vulnerability disclosure and notification may get manual or automatic responses from various communication systems in different languages. In most cases, we have seen English used in the responses; other languages are used as support (Durumeric et al., 2014; Li et al., 2016; Chen et al., 2024).
  • A finder should be aware that messages in different languages may increase the processing time and cause confusion, especially if disclosure policies are presented in non-native languages to a finder or a receiving party.
  • A finder may need to put extra effort into reformulating the messages based on the limitations of the disclosure form or editors in the provided text or system when using stakeholders’ communication systems (Chen et al., 2024).
To inform the intended disclosure or notification, it’s important to provide the motivation, mitigation scheduling, and legal terms, if possible. As we observed in the selected publications, disclosure policies and equivalent documentation are not widely implemented. This has resulted in both finders and receiving parties having inefficient communication.
@@
  • A finder should be aware of the stance of stakeholders during disclosure or notification. Not every stakeholder may want to adopt the best practices for different reasons. This has been reported in two disclosure operations (Vanhoef and Piessens, 2017a; Van Hove et al., 2023) with vendors refusing to participate in the remediation schedule despite having a national CSIRT as coordinator. Still, we encourage stakeholders to adopt best practices and participate in disclosure or notification operations.
  • We recommend that a finder and support organisations establish an outgoing disclosure policy (van der Ham et al., 2023; Reidsma et al., 2023; National Cyber Security Centre of The Netherlands (2018), NCSC-NL), which includes legal terms, disclosure schedules, message templates, or exemptions, as implemented in (Stivala et al., 2024; Van Hove et al., 2023; Chen et al., 2024). This allows recipients to understand the disclosure procedure from a finder and protects the finder from unwanted behaviours, such as legal actions (Macnish and van der Ham, 2020) and public criticism (Van Hove et al., 2023).
  • We recommend that the stakeholders include a disclosure policy(CERT/CC, 2024; National Cyber Security Centre of The Netherlands (2018), NCSC-NL; van der Ham et al., 2023) or security.txt(Foudil and Shafranovich, 2022) in their disclosure programs or responses. This helps a finder to initiate the disclosure and notification operations with the stakeholders with better preparation.
@@ 5.5. Post-Review
With the fast-changing nature of the network landscape and the growing number of vulnerabilities, it is important to have best practices up to date and mitigate security issues in time. We have seen publications focusing on the exploits, attacks, and network traffic before and after the public disclosure. However, documentation focusing on large-scale vulnerability disclosure and notification is relatively scarce in certain communities, as noted in (Chen et al., 2024; Moura and Heidemann, 2023; Stock et al., 2018b). Nevertheless, we have seen that finders in (Chen et al., 2024; Stivala et al., 2024; Beitis and Vanhoef, [n. d.]) could learn from prior best practices and contribute to large-scale notification. Handling disclosure or notification procedures can be a nerve-racking trial, as revealed in nearly all selected publications. It is crucial to have evolving best practices to help a finder prepare for challenges and ease the stress during operations.
  • We recommend that security researchers, ethical hackers, practitioners, and more finders document their disclosure or notification operations with their experiences, considerations, and outcomes.
  • We recommend that finders review the impact before and after vulnerability disclosure or notification, which can help finders reflect on their progress and improve best practices.
  • A finder can record the challenges and stress encountered during a disclosure or notification operation. The tradeoffs and considerations in different scenarios are also worth recording.
  • A finder can provide a timeline as a figure, like in (Boucher and Anderson, 2022; Van Hove et al., 2023; Moura and Heidemann, 2023; Durumeric et al., 2014; Stock et al., 2018b, 2016a) or as text, like in (Chen et al., 2024; Stivala et al., 2024), for each disclosure or notification stage, which can help understand the operation development and its impact over time.
  • A finder can track the number of remaining vulnerable systems before and after disclosure or notification, if the vulnerable systems can be traced through network scanning, stakeholder engagement, or user communication.
  • A finder can track the mitigation progress from weeks (Durumeric et al., 2014), months (Çetin et al., 2019b; Chen et al., 2024), or more than a year (Stivala et al., 2024) if the situation permits. The tracking update can be presented as a webpage (Vanhoef and Piessens, 2017a; Boucher and Anderson, [n. d.]) , experience reports (Moura and Heidemann, 2023), or follow-up academic publications (Durumeric et al., 2014; Stivala et al., 2024; Boucher and Anderson, 2022; Stock et al., 2018b; Vanhoef and Piessens, 2018b) to present disclosure or notification updates.
Last but not least, as we have observed in the academic security research community, academic researchers as finders may show different preferences and limitations compared to bug bounty hunters or vendors as finders. Researchers may need to handle the publication cycle aside from the disclosure or notification procedure, which brings extra time constraints and stress in their work. Further, we have observed that not every academic researcher has the capacity and prior experience to conduct disclosure or notification operations and finish the documentation within the academic publication period. We believe such a situation can arise in different forms for finders in some other communities. With the increasing adoption of support mechanisms such as the disclosure and bug bounty program, the stress of conducting vulnerability notification at scale can be mitigated, but not greatly reduced, as discussed in Section 4.2.5. This is the main motivation for proposing best practices with suggestions in each operation stage to help finders, considering scalability, effective communication, and finder protection.
We have answered our final research question presented in Section 1.1. However, more perspectives on vulnerability disclosure and notification from different finders and stakeholders also require attention. We hope more finders and other stakeholders can benefit from our recommendations and contribute to the security communities by documenting their own experiences and improving best practices.